# Evaluating Autoencoders for Parametric and Invertible Multidimensional Projections

### v2

@@ 1 Introduction
Multidimensional projections, also known as dimensionality reduction (DR) methods, are well-established and frequently used to analyze high-dimensional data visually [Jol86]. DR methods reduce high-dimensional data to a lower-dimensional projection, usually to 2D or 3D, while trying to preserve relationships, i.e., distances and neighborhoods [NA19]. A shortcoming of these approaches is that the result can be hard to interpret since the relationships are projected non-linearly. To address this, researchers proposed to enhance the visualization through layout enrichments, showing the distortions and gradients between projected points [NA19, EAS*21, DMKE24]. However, we argue that these methods benefit from projections that are parametric and invertible. (1) Parametric methods use a model with learnable parameters to control the mapping between high- and low-dimensional spaces. Doing so allows the projection of new data points, real and synthetic, without recalculating all pairwise relationships. This is faster and keeps the projection stable [SMG21]. (2) Invertible methods have a smooth mapping from the projection space back to the original high-dimensional space, allowing us to generate new data from any arbitrary position of the projection [vWvO03], e.g., for interactive counterfactual generation [SRK24]. However, DR methods, such as t-SNE [vdMaa09], are generally not parametric or invertible. In recent years, neural networks (NNs) have been used to learn mappings from high-dimensional space to the projection space [EHT20], and vice versa [EAS*21, HHKS23]. However, while these approaches can be applied to arbitrary projections, they only consider each mapping direction individually. In this work, we explore and evaluate different autoencoders (AEs) [BKG23] to support parametric mapping and inversion jointly through their encoder-decoder architecture. More specifically, we compare the individual feed-forward NNs proposed by Espadoto et al. [EAS*21] and Hinterreiter et al. [AEC*22] with standard AEs and variational AEs with custom loss functions for representing the projection space in the latent space. We measure their ability to parametrically project and inverse project data points and compare them qualitatively by visually assessing their outputs. Additionally, we analyze the smoothness of the projection through gradient maps. Overall, we contribute the following:
@@
    We performed an evaluation comparing the three architectures quantitatively and qualitatively on four datasets using t-SNE.
@@ 2 Related Work on Multidimensional Projections
Let $D=\{x_{i}\}_{1\leq i\leq n}$ be a high-dimensional dataset with $d$ dimension and $n$ samples $x_{i}\in\mathbb{R}^{d}$. A projection method $P$ maps $D$ to $P(D)=\{P(x_{i})|x_{i}\in D\}=\{y_{i}\}_{1\leq i\leq n}$, where $P(D)\subset\mathbb{R}^{q}$ with $q\ll d$. For our case, $q=2$; thus, $P(D)$ can be visualized in a 2-dimensional scatterplot. Projection methods have been extensively studied and assessed in several surveys [Yin07, CG15, EMK*19, NA19]. They can be classified into linear and non-linear methods [LMW*17, EMK*19]. Additionally, they either preserve global structure or local neighborhoods. Linear projection methods like PCA [Jol86] can be computed very efficiently and preserve the global structure of the data. MDS [Kru78] is an example of a global, non-linear projection method. There exist many non-linear methods that put stronger weight on preserving local neighborhood structures at the expense of global structure fidelity. Such methods include t-SNE [vdMH08] or UMAP [MHM18]. Autoencoder-based approaches generally also fall into this category [WYZ16].
@@
Parametric Projections Methods: Standard non-linear DR methods [Kru78, vdMH08, MHM18] are non-parametric, needing complete recalculation when projecting new data points [HHKS23]. Parametric approaches [BBH12] address this limitation, e.g., by using NNs to learn mapping functions into lower-dimensional space [HS06, vdMaa09]. Van der Maaten [vdMaa09] used a feed-forward NN to build parametric t-SNE. Similarly, parametric UMAP [SMG21] uses NNs, including autoencoders, designed to replace the non-parametric embedding step. By training NNs to infer 2D coordinates for input domain data points, Espadoto et al. [EHT20] showed that NNs with sufficient size can effectively approximate many existing non-parametric methods. HyperNP [AEC*22] uses NNs to approximate projection techniques across hyperparameters (e.g., perplexity for t-SNE), allowing for their interactive exploration. ParaDime [HHKS23] streamlines the creation of NN-based DR approaches by proposing a grammar specific to parametric DR.
@@
Inverse Projections: Inverse projections are functions $P^{-1}:\mathbb{R}^{q}\rightarrow\mathbb{R}^{d}$ which are smooth and minimize the cost $\sum_{x\in D}\|P^{-1}(P(x))-x\|$ with projection $P$ and $\|\cdot\|$ denoting the $L_{2}$ norm. Only a handful of projections are inherently inversely defined, e.g., UMAP [MHM18]. An inverse mapping is not explicitly available for most other projection methods, requiring alternative methods, such as NNs [EAS*21]. One of the early methods introduced a global interpolation-based technique to generate new data points [vWvO03]. Later, iLAMP [dSBI*12], which builds on LAMP [JCC*11], addresses inverse projection by focusing on preserving local relationships. Amorim et al. [ABM*15] later refined this by using a radial basis function kernel for interpolation. Blumberg et al. [BWT*24] invert MDS projections by leveraging geometrical relationships between data points through multilateration. Recently, deep learning was employed for creating inverse projections. The feed-forward NNs proposed by Espadoto et al. [EAS*21] learn the mapping from the projection space back to the high-dimensional space.
@@ 3 Training Autoencoders for Multidimensional Projections
We train AEs to jointly create a parametric projection $P$ and inverse projection $P^{-1}$. AEs are a class of NNs designed for unsupervised learning of efficient data representations [BKG23]. AEs were already well-established in the context of DR [HS06]. However, a standard AE will learn a latent space representation optimized for data compression and feature extraction. Thus, modified loss functions were proposed to impose a structure on the latent space of an AE, e.g., SSNP [EHT21] and ShaRP [MTB24] integrate pseudo-labels from clustering into their loss functions. However, these approaches do not allow for inverting arbitrary user-defined projections.
@@
An AE consists of an encoder and a decoder (Sec. 3). The encoder $\mathrm{Enc}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{q}$ learns a mapping from the high-dimensional input space into a lower-dimensional latent space $\hat{Y}=\{\hat{y}_{i}\}_{i=1}^{N}$, with $\hat{y}_{i}=\mathrm{Enc}(x_{i})\in\mathbb{R}^{q}$ and NN data points. In contrast, the decoder $\mathrm{Dec}:\mathbb{R}^{q}\rightarrow\mathbb{R}^{d}$ aims to map these latent representations back into the original input space $\hat{x}_{i}=\mathrm{Dec}(y_{i})\in\mathbb{R}^{d}$. A standard AE aims to learn a compressed latent encoding $y_{i}$ while reconstructing the input $x_{i}$ as accurately as possible. During training, we feed the input $x_{i}$ forward through the encoder and decoder to obtain a reconstruction $\hat{x}_{i}$, then compute a reconstruction loss, i.e., end-to-end. A common choice for this loss is the mean squared error (MSE) $\mathcal{L}_{rec}(x_{i},\hat{x}_{i})=MSE(x_{i},\hat{x}_{i}):=\|x_{i}-\hat{x}_{i}\|^{2}$. This error is backpropagated through the AE, and its parameters are updated accordingly. We use modern training methods for deep AEs, including batches of data and stochastic gradient descent [KB15]. We use standard loss functions, like the mean squared error (MSE) and the Kullback-Leibler divergence ($D_{\textrm{KL}}$), a measure of the difference of two probability distributions $A$ and $B$. In the following paragraph, we describe our proposed NN architectures.
@@
Variational Autoencoder with Latent Loss Term (VAEL): Instead of directly predicting the latent variable $\hat{y}$, the encoder predicts the parameters of the normal distribution $\mu$ and $\sigma^{2}$. In training, $\hat{y}$ is sampled from a 2D normal distribution, i.e., $\hat{y}\sim\mathcal{N}(\mu,\sigma^{2})$ with $\hat{y}$, $\mu$, $\sigma^{2}\in\mathbb{R}^{q}$. The VAE is still trained end-to-end, using the reparameterization trick, to backpropagate through the sampling operation [KW14]. For our VAEL, we incorporate the evidence lower bound loss (ELBO) used for VAE training as:
@@
We follow the framework of the $\beta-VAE$ [HMP*17], which adds an additional $\beta$-parameter to balance conformity to the prior normal distribution and reconstruction quality. We sampled various combinations and found that $\alpha=1.0$ and $\beta=0.1$ achieve generally good quality in terms of MSE on test data. The effect of varying $\beta$ is shown in Fig. 2. We discuss the selection of $\alpha$ and $\beta$ in Sec. 5. Similar to Chen et al. [CG24], we use $\mu$ as a 2D coordinate to create parametric projections, i.e., for inference only, since it is the mean and mode of the normal distribution.
@@ 3.1 Data Preprocessing and Training
We derive the topologies of our AEs from the configurations of the NNs proposed by Espadoto et al., and Appleby et al. [EAS*21, AEC*22]. For P&R, we use these topologies directly; for AEL and VAEL, we combine the NNs to create a bottleneck according to our descriptions. Using the Adam optimizer [KB15], we observed that the training of our AEs converged similarly fast. Thus, we set the number of training epochs to 50, allowing all to converge. The learning rate is set to 0.001, and the batch size is set to 32, which are the typical values. We use batch normalization after each layer and dropout regularization with a probability of $0.25$. For all details, refer to the source code. In our experiments, we standardize the input data per dimension, i.e., the dataset and the projection. When projecting or reconstructing data using the encoder or decoder, we apply the inverse of the standardization to the output, enabling us to recover the representation in the original space and the projection space, respectively. This is possible since standardization is an invertible linear transformation due to $\sigma^{2}\in\mathbb{R}^{+}$.
@@ 4 Evaluation
Since there is no ground truth for many points $\hat{y}\in\mathbb{R}^{2}\setminus P(D)$, we use gradient maps [EAS*21] to evaluate the inverse projection for those points. For each $\hat{y}$, we compute a pseudo-derivative of $P^{-1}$ measuring the difference between the horizontal and vertical neighbors of a pixel defined as $G(q)=\sqrt{\|P^{-1}(\hat{y}_{left})-P^{-1}(\hat{y}_{right})\|^{2}+\|P^{-1}(\hat{y}_{up})-P^{-1}(\hat{y}_{down})\|^{2}}$. This allows us to visualize the rate of change in the high-dimensional space and quantitatively assess the inverse projection’s gradients. We compare the results across four datasets (Tab. 1) using t-SNE with standard parameters [vdMH08] as the ground-truth projection to be learned.
@@
Rings [BWT*24] 33 33 $180$ 0.0% Synthetic
HAR [AGO*12] $561$ $120$ $735$ 0.0% Sensor Data
MNIST [LCB98] $784$ $330$ $70000$ 82.9% Images
Fashion MNIST [XRV17] $784$ $187$ $3000$ 50.2% Images
Table 1: Evaluation datasets with dimensionality $\mathbf{d}$, intrinsic dimensionality $\rho_{\mathbf{d}}$, dataset size $\mathbf{n}$, and sparsity $\gamma_{\mathbf{n}}$ [EMK*19].
@@ 4.1 Quantitative Comparison
The average MSE of the parametric projection (Tab. 2, top) shows how well the different methods project high-dimensional data points for a given projection method $P$, i.e., t-SNE. The P&R applied to the Rings dataset results in a high average MSE, suggesting that the projection performs the worst comparatively to the others, while AEL outperformed the other two. For high-dimensional datasets (i.e., HAR, MNIST, and Fashion MNIST), AEL and VAEL cannot outperform P&R. VAEL outperformed AEL in all benchmarks except HAR. The average MSE of the inverse projection (Tab. 2, bottom) shows how well the approaches can create a high-dimensional data point from a 2D point in the projection. AEL outperformed the other approaches for the Rings and HAR datasets. For Fashion MNIST, it performed the worst. Comparing P&R with VAEL shows that P&R generally outperformed VAEL except for the low-dimensional Rings dataset. The average gradient (Fig. 3, bottom right) measures the average rate of change of the high-dimensional inverse point when moving to a neighboring pixel in the 2D projection. The average gradient increases with the intrinsic dimensionality of the dataset and generally decreases from R&P to VAEL, suggesting that the VAEL generally produces smoother gradient maps. Training times generally increase with the intrinsic dimensionality and size of the dataset (see supplementary material). R&P requires the longest time to train, mainly because it involves training two separate networks. Additionally, the VAEL takes longer to train than the AEL. Overall, inference times are similar and generally low for all the datasets, staying between 0.01s and 0.04s for all architectures.
@@ 5 Discussion and Future Work
The results for P&R show that a feed-forward NN generally works best for creating a parametric projection. Similarly, P&R outperformed AEL and VAEL for generating inverse projections of MNIST and Fashion MNIST. The architectures of AEL and VEAL are directly derived from P&R. However, AEL and VEAL may benefit from additional and larger layers, requiring a validation set to determine their hyperparameters. Our results suggest that the joint training of $P$ and $P^{-1}$ using AEs generally yields smoother inverse projections. In terms of smoothness, VAEL gives the best results with low average gradients and smaller areas of high gradients (Fig. 3), showing that incorporating the $D_{KL}$ loss term has a smoothing effect. Our evaluation suggests that this effect comes at the cost of the parametric and inverse projection accuracy. However, the smoothing strength is controllable through the parameters of VAEL.

### v3

@@ 1 Introduction
Multidimensional projections, also known as dimensionality reduction (DR) methods, are well-established and frequently used to analyze high-dimensional data visually [19]. DR methods reduce high-dimensional data to a lower-dimensional projection, usually to 2D or 3D, while trying to preserve relationships, i.e., distances and neighborhoods [27]. A shortcoming of these approaches is that the result can be hard to interpret since the relationships are projected non-linearly. To address this, researchers proposed to enhance the visualization through layout enrichments, showing the distortions and gradients between projected points [27, 11, 9]. However, we argue that these methods benefit from projections that are parametric and invertible. (1) Parametric methods use a model with learnable parameters to control the mapping between high- and low-dimensional spaces. Doing so allows the projection of new data points, real and synthetic, without recalculating all pairwise relationships. This is faster and keeps the projection stable [28]. (2) Invertible methods have a smooth mapping from the projection space back to the original high-dimensional space, allowing us to generate new data from any arbitrary position of the projection [32], e.g., for interactive counterfactual generation [29]. However, DR methods, such as t-SNE [31], are generally not parametric or invertible. In recent years, neural networks (NNs) have been used to learn mappings from high-dimensional space to the projection space [13], and vice versa [11, 16]. However, while these approaches can be applied to arbitrary projections, they only consider each mapping direction individually. In this work, we explore and evaluate different autoencoders (AEs) [4] to support parametric mapping and inversion jointly through their encoder-decoder architecture. More specifically, we compare the individual feed-forward NNs proposed by Espadoto et al. [11] and Appleby et al. [3] with standard AEs and variational AEs with custom loss functions for representing the projection space in the latent space. We measure their ability to parametrically project and inverse project data points and compare them qualitatively by visually assessing their outputs. Additionally, we analyze the smoothness of the projection through gradient maps. Overall, we contribute the following:
@@
    We perform an evaluation comparing the three architectures quantitatively and qualitatively on four datasets using t-SNE.
@@ 2 Related Work on Multidimensional Projections
Let $D=\{x_{i}\}_{1\leq i\leq n}$ be a high-dimensional dataset with $d$ dimensions and $n$ samples $x_{i}\in\mathbb{R}^{d}$. A projection method $P$ maps $D$ to $P(D)=\{P(x_{i})|x_{i}\in D\}=\{y_{i}\}_{1\leq i\leq n}$, where $P(D)\subset\mathbb{R}^{q}$ with $q\ll d$. For our case, $q=2$; thus, $P(D)$ can be visualized in a 2-dimensional scatterplot. Projection methods have been extensively studied and assessed in several surveys [35, 8, 14, 27]. They can be classified into linear and non-linear methods [24, 14]. Additionally, they either preserve global structure or local neighborhoods. Linear projection methods like PCA [19] can be computed very efficiently and preserve the global structure of the data. MDS [22] is an example of a global, non-linear projection method. There exist many non-linear methods that put stronger weight on preserving local neighborhood structures at the expense of global structure fidelity. Such methods include t-SNE [30] or UMAP [26]. Autoencoder-based approaches generally also fall into this category [33].
@@
Parametric Projections Methods: Standard non-linear DR methods [22, 30, 26] are non-parametric, needing complete recalculation when projecting new data points [16]. Parametric approaches [6] address this limitation, e.g., by using NNs to learn mapping functions into lower-dimensional space [17, 31]. Van der Maaten [31] used a feed-forward NN to build parametric t-SNE. Similarly, parametric UMAP [28] uses NNs, including autoencoders, designed to replace the non-parametric embedding step. By training NNs to infer 2D coordinates for input domain data points, Espadoto et al. [13] showed that NNs with sufficient size can effectively approximate many existing non-parametric methods. HyperNP [3] uses NNs to approximate projection techniques across hyperparameters (e.g., perplexity for t-SNE), allowing for their interactive exploration. ParaDime [16] streamlines the creation of NN-based DR approaches by proposing a grammar specific to parametric DR.
@@
Inverse Projections: Inverse projections are functions $P^{-1}:\mathbb{R}^{q}\rightarrow\mathbb{R}^{d}$ which are smooth and minimize the cost $\sum_{x\in D}\|P^{-1}(P(x))-x\|$ with projection $P$ and $\|\cdot\|$ denoting the $L_{2}$ norm. Only a handful of projections are inherently inversely defined, e.g., UMAP [26]. An inverse mapping is not explicitly available for most other projection methods, requiring alternative methods, such as NNs [11]. One of the early methods introduced a global interpolation-based technique to generate new data points [32]. Later, iLAMP [10], which builds on LAMP [18], addresses inverse projection by focusing on preserving local relationships. Amorim et al. [1] later refined this by using a radial basis function kernel for interpolation. Blumberg et al. [5] invert MDS projections by leveraging geometrical relationships between data points through multilateration. Recently, deep learning was employed for creating inverse projections. The feed-forward NNs proposed by Espadoto et al. [11] learn the mapping from the projection space back to the high-dimensional space.
@@ 3 Training Autoencoders for Multidimensional Projections
We train AEs to jointly create a parametric projection $P$ and inverse projection $P^{-1}$. AEs are a class of NNs designed for unsupervised learning of efficient data representations [4]. AEs were already well-established in the context of DR [17]. However, a standard AE will learn a latent space representation optimized for data compression and feature extraction. Thus, modified loss functions were proposed to impose a structure on the latent space of an AE, e.g., SSNP [12] and ShaRP [25] integrate pseudo-labels from clustering into their loss functions. However, these approaches do not allow for inverting arbitrary user-defined projections.
@@
An AE consists of an encoder and a decoder (Sec. 3). The encoder $\mathrm{Enc}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{q}$ learns a mapping from the high-dimensional input space into a lower-dimensional latent space $\hat{Y}=\{\hat{y}_{i}\}_{i=1}^{n}$, with $\hat{y}_{i}=\mathrm{Enc}(x_{i})\in\mathbb{R}^{q}$ and $n$ data points. In contrast, the decoder $\mathrm{Dec}:\mathbb{R}^{q}\rightarrow\mathbb{R}^{d}$ aims to map these latent representations back into the original input space $\hat{x}_{i}=\mathrm{Dec}(y_{i})\in\mathbb{R}^{d}$. A standard AE aims to learn a compressed latent encoding $y_{i}$ while reconstructing the input $x_{i}$ as accurately as possible. During training, we feed the input $x_{i}$ forward through the encoder and decoder to obtain a reconstruction $\hat{x}_{i}$, then compute a reconstruction loss, i.e., end-to-end. A common choice for this loss is the mean squared error (MSE) $\mathcal{L}_{rec}(x_{i},\hat{x}_{i})=MSE(x_{i},\hat{x}_{i}):=\|x_{i}-\hat{x}_{i}\|^{2}$. This error is backpropagated through the AE, and its parameters are updated accordingly. We use modern training methods for deep AEs, including batches of data and stochastic gradient descent [20]. We use standard loss functions, like the mean squared error (MSE) and the Kullback-Leibler divergence ($D_{\textrm{KL}}$), a measure of the difference of two probability distributions $A$ and $B$. In the following paragraph, we describe our proposed NN architectures.
@@
Variational Autoencoder with Latent Loss Term (VAEL): Instead of directly predicting the latent variable $\hat{y}$, the encoder predicts the parameters of the normal distribution $\mu$ and $\sigma^{2}$. In training, $\hat{y}$ is sampled from a 2D normal distribution, i.e., $\hat{y}\sim\mathcal{N}(\mu,\sigma^{2})$ with $\hat{y}$, $\mu$, $\sigma^{2}\in\mathbb{R}^{q}$. The VAE is still trained end-to-end, using the reparameterization trick, to backpropagate through the sampling operation [21]. For our VAEL, we incorporate the evidence lower bound loss (ELBO) used for VAE training as:
@@
We follow the framework of the $\beta-VAE$ [15], which adds an additional $\beta$-parameter to balance conformity to the prior normal distribution and reconstruction quality. We sampled various combinations and found that $\alpha=1.0$ and $\beta=0.1$ achieve generally good quality in terms of MSE on test data. The effect of varying $\beta$ is shown in Fig. 2. We discuss the selection of $\alpha$ and $\beta$ in Sec. 5. Similar to Chen et al. [7], we use $\mu$ as a 2D coordinate to create parametric projections, i.e., for inference only, since it is the mean and mode of the normal distribution.
@@ 3.1 Data Preprocessing and Training
We derive the topologies of our AEs from the configurations of the NNs proposed by Espadoto et al., and Appleby et al. [11, 3]. For P&R, we use these topologies directly; for AEL and VAEL, we combine the NNs to create a bottleneck according to our descriptions. Using the Adam optimizer [20], we observed that the training of our AEs converged similarly fast. Thus, we set the number of training epochs to 50, allowing all to converge. The learning rate is set to 0.001, and the batch size is set to 32, which are the typical values. We use batch normalization after each layer and dropout regularization with a probability of $0.25$. For all details, refer to the source code. In our experiments, we standardize the input data per dimension, i.e., the dataset and the projection. When projecting or reconstructing data using the encoder or decoder, we apply the inverse of the standardization to the output, enabling us to recover the representation in the original space and the projection space, respectively. This is possible since standardization is an invertible linear transformation due to $\sigma^{2}\in\mathbb{R}^{+}$.
@@ 4 Evaluation
Since there is no ground truth for many points $\hat{y}\in\mathbb{R}^{2}\setminus P(D)$, we use gradient maps [11] to evaluate the inverse projection for those points. For each $\hat{y}$, we compute a pseudo-derivative of $P^{-1}$ measuring the difference between the horizontal and vertical neighbors of a pixel defined as $G(\hat{y})=\sqrt{\|P^{-1}(\hat{y}_{left})-P^{-1}(\hat{y}_{right})\|^{2}+\|P^{-1}(\hat{y}_{up})-P^{-1}(\hat{y}_{down})\|^{2}}$. This allows us to visualize the rate of change in the high-dimensional space and quantitatively assess the inverse projection’s gradients. We compare the results across four datasets (Tab. 1) using t-SNE with standard parameters [30] as the ground-truth projection to be learned.
@@
Rings [5] 33 33 $180$ 0.0% Synthetic
HAR [2] $561$ $120$ $10299$ 0.0% Sensor Data
MNIST [23] $784$ $330$ $70000$ 82.9% Images
Fashion MNIST [34] $784$ $187$ $70000$ 50.2% Images
Table 1: Evaluation datasets with dimensionality $\mathbf{d}$, intrinsic dimensionality $\rho_{\mathbf{d}}$, dataset size $\mathbf{n}$, and sparsity $\gamma_{\mathbf{n}}$ [14].
@@ 4.1 Quantitative Comparison
The average MSE of the parametric projection (Tab. 2, top) shows how well the different methods project high-dimensional data points for a given projection method $P$, i.e., t-SNE. The P&R applied to the Rings dataset results in a high average MSE, suggesting that the projection performs the worst comparatively to the others, while AEL outperformed the other two. For high-dimensional datasets (i.e., HAR, MNIST, and Fashion MNIST), AEL and VAEL cannot outperform P&R. VAEL outperformed AEL in the MNIST and Fashion MNIST benchmarks. The average MSE of the inverse projection (Tab. 2, bottom) shows how well the approaches can create a high-dimensional data point from a 2D point in the projection. AEL outperformed the other approaches for the Rings and HAR datasets. For Fashion MNIST, VAEL performed the worst. Comparing P&R with VAEL shows that P&R generally outperformed VAEL except for the low-dimensional Rings dataset. The average gradient (Fig. 3, bottom right) measures the average rate of change of the high-dimensional inverse point when moving to a neighboring pixel in the 2D projection. The average gradient increases with the intrinsic dimensionality of the dataset and generally decreases from P&R to VAEL, suggesting that the VAEL generally produces smoother gradient maps. Training times generally increase with the intrinsic dimensionality and size of the dataset (see supplementary material). P&R requires the longest time to train, mainly because it involves training two separate networks. Additionally, the VAEL takes longer to train than the AEL. Overall, inference times are similar and generally low for all the datasets, staying between 0.01s and 0.04s for all architectures.
@@ 5 Discussion and Future Work
The results for P&R show that a feed-forward NN generally works best for creating a parametric projection. Similarly, P&R outperformed AEL and VAEL for generating inverse projections of MNIST and Fashion MNIST. The architectures of AEL and VAEL are directly derived from P&R. However, AEL and VAEL may benefit from additional and larger layers, requiring a validation set to determine their hyperparameters. Our results suggest that the joint training of $P$ and $P^{-1}$ using AEs generally yields smoother inverse projections. In terms of smoothness, VAEL gives the best results with low average gradients and smaller areas of high gradients (Fig. 3), showing that incorporating the $D_{KL}$ loss term has a smoothing effect. Our evaluation suggests that this effect comes at the cost of the parametric and inverse projection accuracy. However, the smoothing strength is controllable through the parameters of VAEL.